On March 31st, 2017 Alexander Corporation sold equipment for cash of $25,000. The equipment had been purchased on June 1st, 2015 for $42,000 and at that time they had estimated a $15,000 residual value at the end of the equipment's 4-year useful life. Alexander Corporation has a December 31st year end, adjusts its records annually, and uses the double declining/diminishing balance method of depreciation for this equipment.

Prepare the appropriate journal entrie(s) to record the disposal of the equipment on
March 31st, 2017. Please show all calculations to receive full marks. Round all amounts to the nearest dollar.
Answer #1

1. Cost of equipment purchased on 01.06.2015 is $42,000

2. Estimated residual value at the end of 4 years is $15,000

3. Useful life of the asset is 4 years

4. Depreciation rate = 1/useful life * 100 = (1/4) * 100 = 25%

5. Double-declining balance formula = 2 * cost of equipment * depreciation rate

6. Depreciation for year 1, 2015 is = (2 * $42,000 * 25%)/2 = $10,500

7. Depreciation for year 2, 2016 is = 2 * ($42,000 - $10,500) * 25% = $15,750

8. Carrying value of equipment at the end of year 2 i.e. 2016 = $42,000 - $10,500 - $15,750 = $15,750

9. Depreciation for year 3, 31st March, 2017 is = (2 * $15,750 * 25%)/2 = $3,938

Since the estimated residual value at the end of useful life is $15,000, the depreciation for 31st March, 2017 is restricted to $750.

10. Accumulated depreciation of equipment till 31st march, 2017 = $10,500 + $15,750 + $750 = $27,000

11. Carrying value of equipment at 31st March, 2017 = $42,000 - $10,500 - $15,750 - $750 = $15,000

12. Sale value of the equipment at 31st March, 2017 = $25,000

Journal entry on disposal of equipment on 31st March, 2017:

Particulars Debit (Dr.) Credit (Cr.)

Cash A/c ................... Dr. $25,000

Accumulated depreciation A/c .................... Dr. $27,000

To Equipment A/c ..................................... Cr. $42,000

To Gain on disposal of Equipment A/c ........Cr. $10,000